Find two numbers that have a sum of 60 and a difference of 36
Answer by ankor@dixie-net.com(22740)

Find two numbers that have a sum of 60 and a difference of 36
a + b = 60
a - b = 36
-------------addition eliminates b find a
2a = 96
a = 96/2
a = 48
:
find b
48 + b = 60
b = 60 - 48
b = 12
;
:
Check: 48 - 12 = 36